# Missing Money and Market-Based Adequacy in Deeply Decarbonized Power Systems with Long-Duration Energy Storage

### v1

@@ III Case study setup
We study a diverse set of near-zero emission systems, which we identify by varying annual EI limits from 15 to 0.1 gCO2/kWh [^7] The selected EIs align with recent Future Energy Scenarios [39]. The model includes emission caps, which are calculated as the product of EI and the expected annual demand ($\sum_{t,\omega}[w_{\omega,t}\delta_{\omega}(D_{\omega,t}^{fix}+D_{\omega,t}^{flex})]$). Assuming 90% CCS efficiency, CCGT with CCS emits 37.8 gCO2/kWh. LDES is excluded from the optimal mix without an emission cap, concurring with previous research on the need for technological advancements [5].
We construct the demand curves for the CM as follows. Following GB CM guidance [40], the curve has two downward-slopping segments, and their breakpoints are: (0.965•$CT$, 1.5•net-CONE), ($CT$, net-CONE), (1.035•$CT$, 0). The target capacity $CT$ is established as defined in II-C.
@@
FOM costs per MWh of storage energy capacity.
@@ IV-A The missing money assessment
We derive a benchmark mix from EOM VOLL simulations for each EI case. Fig. 1 shows the capacity mix and LDES duration. Tighter EI targets reduce CCGT-CCS capacity while increasing wind, solar, and LDES. Batteries play a minor, auxiliary role [^8] We do not model ancillary services, where batteries are expected to be essential.. Despite technological changes, non-VRE capacity remains near 80% of peak demand, as firm capacity needs slowly decline with VRE expansion. Fig. 1b concurs with [3], with LDES duration rising sharply as decarbonisation deepens.
Figure 1: Capacity results for the ‘EOM with VOLL’ simulations. a) shows the installed power capacity as a percentage of peak demand, while b) shows the installed duration of LDES.
Fixing capacities at the levels from Fig. 1, we dispatch resources under a PC to estimate how much money each technology is missing to break even. Fig. 2 shows net revenues from the energy market per technology for two sets of time steps that have energy prices Below and Above or at PC, respectively. For storage, we estimate the cost of energy discharged based on the average charging cost of stored energy at the time of discharge. Appendix 1 describes its calculation.
CCGT-CCS recovers more than 80% of its costs during intervals with prices above the PC in all EI cases. In contrast, storage recovers between 30–80% in these intervals, with this fraction decreasing as emission limits tighten. This trend translates into a diminishing missing money problem for storage. Lower charging prices, driven by higher VRE generation and curtailment, result in reduced storage costs. At the same time, rising carbon costs increase the SRMC of CCGT-CCS, raising discharging prices when CCGT-CCS is the marginal producer, and thereby widening arbitrage spreads. We also observe that, unlike systems without storage, the introduction of the PC not only reduces the Above or at PC net revenues, but also affects net revenues during hours when prices are below the PC, highlighting the intertemporal nature of the missing money problem recently emphasized by [41]. VRE technologies recover a large share of their costs in all cases and are less exposed to missing money. In contrast, CCGT-CCS remains the most vulnerable, as caps suppress scarcity rents that would otherwise support cost recovery.
@@ IV-B Capacity credit estimates
We estimate the marginal CCs presented in Fig. 1 following Algorithm 1. For storage, we compute up to 56 marginal CCs, [^9] For less stringent EI limits, fewer point estimates are required, as a CC of 1 is reached at shorter durations. with one CC per ES duration. We fit a piecewise linear curve with four segments to the CC estimates. The curve closely follows the estimates, yielding an $R^{2}$ value exceeding 0.99 in all cases, which confirms the appropriateness of the piecewise linear approach. Fig. 3 illustrates a critical dynamic: marginal CC estimates are highly sensitive to the underlying resource mix. In general, storage requires longer durations to achieve full accreditation as EI limits become more stringent and dispatchable capacity diminishes. The marginal CC for batteries is higher than for LDES for equivalent durations, owing to better operational characteristics, primarily higher charging efficiency, facilitating greater utilisation under narrower charging/discharging price spreads. We observe that the curves for the 1 and 0.1gCO2/kWh cases are surprisingly close and even cross in the Battery case. To interpret and relate the results on CC with the missing money problem, we contrast the estimated CC to the required CC for all technologies in Fig. 4a, where we also present the net CONEs per marginally accredited MW.
@@
Figure 3: Marginal capacity credit estimates as a function of storage duration. Markers indicate credit obtained with duration from EOM VOLL.
For all cases except 1gCO2/kWh, we observe that the estimated credits are close to the required values and that net CONEs are similar among technologies. To explain the differences in the 1gCO2/kWh case, we examine how ES influences the estimated CC under three paradigms for storage dispatch in Step 2 of algorithm 1. In the first paradigm, labelled unconstrained in Fig. 4b, the dispatch (model M) decides storage charging and discharging. This is the assumption that we have used so far in the article to estimate CC. In the second paradigm, labelled existing ES charging fixed in Fig. 4b, the dispatch decides storage discharging and keeps the ES charging at levels identified by the SW-maximising problem. Last, in the third paradigm, labelled existing ES charging and discharging fixed in Fig. 4b, the dispatch keeps both charging and discharging variables of ES at the levels identified by the SW-maximising solution.
@@
Figure 4: a) Credited net-CONEs and capacity credits for technologies identified by ‘EOM with VOLL’ simulations. Note for storage, the optimal duration is considered. b) CCGT-CCS marginal 0.01MW unit contribution to EUE reduction under alternative setups.
The difference in EUE reduction (output of Step 2 in Algorithm 1) between the second and the third paradigm is expected. Zachary et al. [15] have shown that when a marginal amount of a generator is added to the mix $\hat{c}_{r}$, the reduction in EUE is greater when the mix already contains storage. This is because once the generator addresses issues in hours that storage was previously discharging, storage discharging can be shifted to other hours with a shortfall of energy.
However, the difference in EUE reduction between the first and second paradigms has not been widely discussed before. That difference is more pronounced in the 1gCO2/kWh case. In that case, the CCGT-CCS capacity is so low that it sometimes generates at full capacity to charge storage; yet even then, storage does not always reach the energy level required to address energy shortfalls in time. Hence, when marginal CCGT-CCS is added, part of the EUE reduction should be attributed to storage discharging due to higher levels of stored energy. This dynamic explains why the estimated credit for storage is lower than its required credit. The same phenomenon explains why VRE’s estimated credit is higher than its required credit, as it partially captures storage’s contributions.
The differences between the first two paradigms are negligible in other cases. In the 15 and 5 gCO2/kWh cases, the CCGT-CCS capacity is relatively high compared to ES, and as such, it is not generating at full capacity when it charges storage. In the 0.1gCO2/kWh case, the emission cap is so low that the effective marginal cost of the CCGT-CCS is greater than $\text{PC}\cdot\eta_{s}^{\text{ch}}\cdot\eta_{s}^{\text{dis}}$.
@@ IV-C Effectiveness of energy plus capacity market
Under a price cap, SW losses are caused by two factors. First, unserved demand is inefficiently distributed among consumers due to a lack of price signals between the PC and VOLL. Second, distorted price signals result in inefficient ES dispatch. In Table II, we show SW losses caused by the second factor for the EOM PC and E+CM. For the benchmark capacity mix dispatched under price cap, we present SW losses for two cases: when both factors are present (EOM PC opt. mix. ineff. dist.) and when only the second factor affects results (EOM PC opt. mix.). Under higher emission limits, welfare losses almost disappear when demand rationing is perfect, indicating that inefficient distribution of unserved energy is the primary reason for SW losses. However, for the 0.1gCO2/kWh, the welfare loss is high even under perfect demand rationing, indicating a substantial impact of PC on storage operations via reduced price spreads.
TABLE II: Social Welfare and Underved Energy for All Runs
EI USE [% of total demand] Welfare Loss [% TC in EOM VOLL]
EOM
VOLL
EOM PC
opt. mix
E+CM
EOM
PC
EOM PC
opt. mix.
ineff. dist.
EOM PC
opt. mix
E+CM
EOM
PC
15.0 0.0043 0.0043 0.0044 0.0143 -0.0140 -0.0003 -0.0055 -0.8536
5.0 0.0034 0.0034 0.0036 0.0133 -0.0521 -0.0078 -0.0158 -0.7101
1.0 0.0018 0.0019 0.0029 0.0088 -0.0338 -0.0114 -0.0675 -0.4253
0.1 0.0027 0.0080 0.0080 0.0294 -0.7876 -0.6862 -0.3773 -1.6923
@@
In order to make the results comparable, for the EOM VOLL case we measure EUE with relation to the fixed part of the true WTP and the ’cut off’ part of the flexible demand above the PC.
@@
Under the E+CM design, we observe welfare losses, but they are significantly lower than those observed with PC but no CRM (EOM PC). For all cases with EI higher than 0.1gCO2/kWh, welfare losses are minimal (below 0.1% of total cost) under the E+CM design, yet growing with a tighter emission limit. However, in the 0.1gCO2/kWh case, the welfare loss is almost 0.4% of total cost under the E+CM design, which might seem high, but it is still roughly half the loss caused by the PC alone (EOM PC opt. mix.). In E+CM design, players account for PC distortions, and their optimal solutions differ from the benchmark mix, primarily in terms of reduced LDES capacity (1.3 MW/66 MWh). That mix is more efficiently dispatched, and the SW is closer to the optimal level under the E+CM compared to a design that supports the benchmark capacity mix. At 1gCO2/kWh, losses are primarily due to misaligned credits, which result in a misaligned mix. In terms of EUE, comparing EOM VOLL and EOM PC opt. mix., we see that the impact of the PC is substantial under the lowest EUE scenario. EUE levels under E+CM closely match the EOM PC opt. mix. levels, except for the 1gCO2/kWh case, which again derives from misaligned credits.

### v2

@@ III Case study setup
We study a diverse set of near-zero emission systems, which we identify by varying annual EI limits from 15 to 0.1 gCO2/kWh [^8] All EI scenarios except 0.1 gCO2/kWh align with recent Future Energy Scenarios [eso_future_2024]. The 0.1 gCO2/kWh case is included primarily to illustrate systems with an even higher reliance on long-duration energy storage.. The model includes emission caps, which are calculated as the product of EI and the expected annual demand ($\sum_{t\omega}[w_{\omega t}\delta_{\omega}(D_{\omega t}^{\text{fix}}+D_{\omega t}^{\text{flex}})]$). Assuming 90% CCS efficiency, CCGT with CCS emits 37.8 gCO2/kWh. LDES is excluded from the optimal mix without an emission cap, concurring with previous research on the need for technological advancements [sepulveda_design_2021].
[…]
@@
FOM costs per MWh of storage energy capacity; WACC is Weighted Average Cost of Capital (WACC) used to annualize capital costs.
@@ IV-A The missing money assessment
We derive a benchmark mix from EOM VOLL simulations for each EI case. Fig. 1 shows the capacity mix and LDES duration. Tighter EI targets reduce CCGT-CCS capacity while increasing wind, solar, and LDES. Batteries play a minor, auxiliary role [^9] We do not model ancillary services, where batteries are expected to be essential. Furthermore, the chronological clustering we use may underestimate installed capacity of short-duration storage systems in large-scale case studies [mannhardt_accurately_2025].. Despite technological changes, non-VRE capacity remains near 80% of peak demand, as firm capacity needs slowly decline with VRE expansion. Fig. 1b concurs with [cebulla_how_2018], with LDES duration rising sharply as decarbonisation deepens.
Figure 1: Capacity results for the EOM VOLL simulations. Panel a shows the installed power capacity as a percentage of peak demand, while panel b shows the installed duration of LDES. In the 15 gCO2/kWh case, no battery capacity is part of the welfare maximising mix, which is why the battery bar does not appear in this cluster and does not appear in subsequent results figures.
Fixing capacities at the levels from Fig. 1, we dispatch resources under a price cap to estimate how much money each technology is missing to break even. Fig. 2 shows net revenues from the energy market per technology for two sets of time steps that have energy prices Below and Above or at price cap, respectively. For storage, we estimate the cost of energy discharged based on the average charging cost of stored energy at the time of discharge. Appendix 1 describes its calculation.
CCGT-CCS recovers more than 80% of its costs during intervals with prices above the price cap in all EI cases. In contrast, storage recovers between 30–80% in these intervals, with this fraction decreasing as emission limits tighten. This trend translates into a diminishing missing money problem for storage. Lower charging prices, driven by higher VRE generation and curtailment, result in reduced storage costs. At the same time, rising carbon costs increase the SRMC of CCGT-CCS, raising discharging prices when CCGT-CCS is the marginal producer, and thereby widening arbitrage spreads. We also observe that, unlike systems without storage, the introduction of the price cap not only reduces the Above or at price cap net revenues, but also affects net revenues during hours when prices are below the price cap, highlighting the intertemporal nature of the missing money problem recently emphasized by [visvesvaran_market-based_2025]. VRE technologies recover a large share of their costs in all cases and are less exposed to missing money. In contrast, CCGT-CCS remains the most vulnerable, as caps suppress scarcity rents that would otherwise support cost recovery.
@@ IV-B Capacity credit estimates
We estimate the marginal CCs presented in Fig. 3 following Algorithm 1. For storage, we compute up to 56 marginal CCs, one per ES duration. We fit a four-segment piecewise linear curve to these estimates; the fitted curves closely follow the point estimates, with an $R^{2}$ exceeding 0.99 in all EI cases, confirming the appropriateness of this approach. The estimated CC curves are monotonically increasing and concave.
Fig. 3 illustrates a critical dynamic: marginal CC estimates are highly sensitive to the underlying resource mix. In general, storage requires longer durations to achieve full accreditation as EI limits tighten, and for certain portfolios, full accreditation may not be achievable at all. This is because storage needs to shift larger amounts of energy as dispatchable capacity reduces. As the amount of ‘energy to shift’ increases, the opportunities to charge (during hours with abundant VRE or other economical supply) and discharge (during hours of scarcity) become more chronologically separated, and the series of ‘scarcity events’ (with price at PC) includes events more chronologically close to each other. When substantial time has lapsed from the previous scarcity event, system storage would be fully charged at the beginning of a ‘scarcity event’. However, when ‘scarcity events’ start shortly after another event ended, system storage would likely go from empty to empty without an opportunity to fully charge in between. In that case, the system would appear to be charging-constrained [^10] We refer to storage as charging-constrained when its adequacy contribution is limited by insufficient opportunities to recharge between chronologically-close scarcity events.. Under charging-constrained conditions, the technologies that can most effectively shift energy have a higher capacity credit. In our test system, this is the case for batteries, which have a higher round-trip efficiency than LDES. This difference in capacity credits between storage technologies with the same duration is a novel observation in our analysis. It reflects that, as the system relies more on storage for adequacy, round-trip efficiency becomes a more material driver of the adequacy contribution of storage, as it is directly linked with its ability to effectively use opportunities to charge between scarcity events.
@@
Figure 3: Marginal capacity credit estimates as a function of storage duration. Markers indicate credit for duration obtained from EOM VOLL simulations.
The charging-constrained conditions complicate the marginal accreditation calculations in two ways. First, when a marginal storage unit is added to the baseline portfolio, it will reduce EUE during charging-constrained hours if the constraint was due to insufficient charging power capacity, but not if there was insufficient surplus energy available before the next shortage period. Second, when a marginal generator is added to the baseline portfolio, it can reduce EUE not only directly in shortage hours, but also indirectly by creating additional surplus energy between scarcity events that storage can shift into a shortage period.
Fig. 4 illustrates the second mechanism by comparing the marginal EUE reduction from adding a small increment of CCGT-CCS in Step 2 of Algorithm 1 under two dispatch setups: (1) unconstrained storage dispatch; (2) existing ES charging fixed. The latter setup fixes storage charging at the level obtained from the price-capped dispatch of the baseline portfolio. When the EUE reduction is larger under (1), the small increment of CCGT-CCS might be used to charge storage. The difference between the EUE reductions under the two setups is most pronounced in the 1 gCO2/kWh case, indicating more severe charging-constrained conditions. The differences are small in other EI cases. For high EIs, the charging-constrained conditions are less severe. For the 0.1 gCO2/kWh case, the effective marginal cost of CCGT-CCS (carbon-inclusive) is economically unattractive for charging LDES, but still attractive for charging batteries.
Figure 4: CCGT-CCS marginal 0.01MW unit contribution to EUE reduction under alternative setups.
As a result of these two mechanisms, under charging-constrained conditions, the capacity credit estimated with Algorithm 1 is less linked with contributions of marginal technologies during scarcity hours and their missing money. For a generator, the estimated capacity credit also captures the generator’s indirect contribution to reducing EUE, which is unlikely to be associated with missing money. For storage, the estimated capacity credit reflects a lower adequacy contribution due to stricter charging constraints for the marginal addition. Hence, the estimated capacity credits are expected to be lower than required for recovering missing money for storage or generators with smaller indirect contributions than the reference generator; and larger for generators with larger indirect contributions than the reference generator. This is indeed the result we observe in Fig. 5, where we compare the missing money-based credit to the EUE-based credit.
In Fig. 5, the bars depict the corresponding net CONEs per (EUE-)accredited MW. Capacity markets have hitherto operated on the basis that, in equilibrium, the net CONE per accredited MW of installed resources should be identical. However, this is not the case for the baseline portfolio with EUE-based credits for our system. As well as the dynamics described above related to storage being charging constrained, we identify two more mechanisms that could explain the differences.
@@
Figure 5: Credited net-CONEs and capacity credits for technologies identified by EOM VOLL simulations. Note for storage, the optimal duration is considered.
First, demand is assumed to be price-elastic above the price cap. With elastic demand, uncapped scarcity prices may vary between the marginal cost of the marginal generator and VOLL, so the missing money per MWh induced by a price cap is not constant but varies across hours. Therefore, if a resource’s output is correlated with the level of missing money per MWh, the missing-money shortfall per MW of that resource can be relatively greater than its marginal contribution to reducing unserved energy. In our case, the output of storage is more correlated with the level of missing money per MW than the reference generator. Storage can be energy-constrained while the reference generator is not, and it discharges only when prices reach the shadow value of stored energy [forsund_hydropower_2007, junge_energy_2022]. If this shadow value lies above the price cap, there will be intermediate scarcity hours in which reference generators incur relatively modest missing-money shortfalls, but storage does not, since it would not have discharged at those times. Prices can also exceed the shadow value of stored energy if storage is power-constrained while discharging, so that these maximal storage discharges are associated with the largest amounts of missing money per MWh. Indeed, in Table II, we observe that the higher the number of intermediate scarcity hours, the larger the differences in net CONE per accredited MW.
Second, some of the missing money for storage is due to the lack of incentives for efficient dispatch in a price-capped environment. In a market without a price cap, storage may optimally charge at prices above $\text{PC}\cdot\eta_{s}^{\text{ch}}\cdot\eta_{s}^{\text{dis}}$ when the discharge value justifies doing so. However, with a price cap, such charging is no longer profitable because the cap truncates the maximum discharge price. Table II confirms that these ‘high-price’ charging events occur in the uncapped dispatch, particularly for LDES under tighter EI targets. By changing these charging decisions and compressing spreads, the price cap can therefore reshape both storage schedules and energy prices outside scarcity hours, further supporting our discussion in Figure 2 that the missing-money distortion has an intertemporal character and is not confined to the scarcity interval. Due to these distorted incentives for dispatch, the baseline portfolio might not maximise the social welfare in an energy-plus-capacity market environment.
These dynamics can be intertwined, and their relative importance depends on the underlying system, so they can shift whether (and how closely) EUE contribution aligns with the missing-money shortfall that a CM is intended to recover. Our results indicate that maintaining this alignment can be challenging in deeply decarbonised systems with high storage penetration and elastic demand, motivating more fundamental derivations of the sufficient conditions under which EFC-based accreditation restores long-run cost recovery, which we leave for future work. Following current industry practice, in the remainder of the results, we therefore proceed using the EUE-based credits to evaluate how the capacity market performs under our case study assumptions.
TABLE II: Average hours per year of specific price and storage-dispatch instances from EOM VOLL case.
Emission intensity (gCO2/kWh)
Parameter 15.0 5.0 1.0 0.1
Intermediate scarcity price region: $\text{PC}<\lambda^{E}_{\omega t}<\text{VOLL}$
Hours 4.93 5.30 15.73 14.28
High price charging hours: $q^{\text{ch}}_{\omega ts}>0$ and $\lambda^{E}_{\omega t}>\text{PC}\cdot\eta_{s}^{\text{ch}}\cdot\eta_{s}^{\text{dis}}$
Battery - 0.63 1.23 0.20
LDES 0.38 1.10 7.50 8.93
@@ IV-C Effectiveness of energy plus capacity market
Under a price cap, SW losses are caused by two factors. First, unserved demand is inefficiently distributed among consumers due to a lack of price signals between the price cap and VOLL. Second, distorted price signals result in inefficient ES dispatch. In Table III, we show SW losses caused by the second factor for the EOM PC and E+CM. For the benchmark capacity mix dispatched under price cap, we present SW losses for two cases: when both factors are present (EOM PC opt. mix. ineff. dist.) and when only the second factor affects results (EOM PC opt. mix.). Under higher emission limits, welfare losses almost disappear when demand rationing is perfect, indicating that inefficient distribution of unserved energy is the primary reason for SW losses. However, for the 0.1gCO2/kWh, the welfare loss is high even under perfect demand rationing, indicating a substantial impact of the price cap on storage operations via reduced price spreads.
TABLE III: Social Welfare and Unserved Energy for All Runs
Welfare Loss [% TC in EOM VOLL]
EI EOM PC opt. mix. ineff. dist. EOM PC opt. mix E+CM EOM PC 15.0 -0.0140 -0.0003 -0.0045 -0.8536 5.0 -0.0521 -0.0078 -0.0156 -0.7101 1.0 -0.0338 -0.0114 -0.0439 -0.4253 0.1 -0.7876 -0.6862 -0.3776 -1.6923
Unserved Energy [% of total demand]
EI EOM VOLL EOM PC opt. mix E+CM EOM PC 15.0 0.0043 0.0043 0.0044 0.0143 5.0 0.0034 0.0034 0.0036 0.0133 1.0 0.0018 0.0019 0.0025 0.0088 0.1 0.0027 0.0080 0.0080 0.0294
@@
For comparability, for the EOM VOLL case we measure EUE relative to the fixed part of the true WTP and the ’cut-off’ part of flexible demand above the PC.
@@
Under the E+CM design, we observe welfare losses, but they are significantly lower than those observed with price cap but no CRM (EOM PC). For all cases with EI higher than 0.1gCO2/kWh, welfare losses are minimal (below 0.1% of total cost) under the E+CM design, yet growing with a tighter emission limit. However, in the 0.1gCO2/kWh case, the welfare loss is almost 0.4% of total cost under the E+CM design, which might seem high, but it is still roughly half the loss caused by the price cap alone (EOM PC opt. mix.). In E+CM design, players account for price cap distortions, and their optimal solutions differ from the benchmark mix, primarily in terms of reduced LDES capacity (-1.3 MW/-60 MWh). That mix is more efficiently dispatched, and the SW is closer to the optimal level under the E+CM compared to a design that supports the benchmark capacity mix. At 1gCO2/kWh, losses are likely driven by the weaker relationship between missing money and EUE contributions. In terms of EUE, comparing EOM VOLL and EOM PC opt. mix., we see that the impact of the price cap is substantial under the lowest EI scenario. EUE levels under E+CM closely match the EOM PC opt. mix. levels, except for the 1gCO2/kWh case, which again derives from weaker relationship between missing money and EUE contributions.
Table IV reports the change in installed capacities in the CM relative to the baseline portfolio. The welfare losses observed in Table III stem from different investment and dispatch incentives in the E+CM vs. the EOM with VOLL case. Under all cases for emission intensity, the E+CM portfolio has higher investment in CCGT-CCS and lower investment in LDES compared to the baseline portfolio. This substitution effect is expected for competing technologies contributing to reliability. Similarly, we observe a substitution between wind and solar investments, with higher EIs having lower wind and higher solar, and lower EIs having higher wind and lower solar than the baseline portfolio. For batteries, we observe mixed results across EIs as they can complement renewables, particularly solar, but also act as a substitute for LDES. We also see that the E+CM mix from 0.1gCO2/kWh case, which, in terms of welfare, performs better than the EOM VOLL mix with price cap, results in lower LDES capacity and higher CCGT-CCS. It indicates that distortions caused by the price cap in the dispatch of storage generally favour conventional generation. Overall, the differences between the baseline portfolio and the E+CM portfolio are relatively small. That’s why we have not included capacity credit surfaces that account for the impact of the portfolio on capacity credits. This small difference is not guaranteed, and that’s why systematic integration of EFC surfaces in a capacity market or resource adequacy process is recommended.
TABLE IV: Absolute Change in Installed Capacities Between E+CM and EOM VOLL Cases
EI PV Wind CCGT CCS Battery* LDES*
15.0 1.05 -0.42 0.41 – / – -0.57 / -23.25
5.0 1.14 -0.39 0.14 0.45 / 2.39 -0.70 / -46.48
1.0 -1.36 2.68 0.06 -0.79 / -2.56 -0.94 / 1.96
0.1 -2.79 0.83 0.20 0.05 / 4.14 -1.32 / -60.66
Absolute changes from EOM VOLL run. Positive values indicate higher capacity in the E+CM case.
For storage technologies, values show the change in power capacity / energy capacity (MW / MWh).